# MoNE: Replacing Redundant Experts with Lightweight Novices for Structured Pruning of MoE

### v1

@@ Abstract
Mixture-of-Experts (MoE) enables efficient scaling of large language models by activating only a subset of experts per input token. However, deploying MoE-based models incurs significant memory overhead due to the need to retain all experts in memory. While structured pruning is promising to reduce memory costs, existing methods often show suboptimal performance and unstable degradation in three dimensions: model architectures, calibration data sources, and calibration sample sizes. This paper proposes Mixture-of-Novices-and-Experts (MoNE), a novel expert pruning method that replaces redundant experts with lightweight novices to achieve effective and robust model compression. MoNE evaluates expert redundancy based on two metrics: access frequency and output variance. Experts exhibiting low usage and stable outputs are pruned and replaced with lightweight novices—unbiased estimations of their original outputs—minimizing performance degradation. Extensive experiments demonstrate that MoNE consistently outperforms baseline methods with minimal accuracy degradation across the three dimensions, confirming its effectiveness and robustness. Notably, it improves the average zero shot accuracy across nine downstream tasks by up to 2.71 under 25% pruning ratio and 3.61 under 50% pruning ratio [^1] The code is available at https://github.com/zxgx/mode-pd..
@@ 1 Introduction
Mixture-of-Experts (MoE) has emerged as a powerful architecture for advancing the capabilities of large language models (LLMs) [20, 21, 25]. MoE-based LLMs achieve higher parameter efficiency than vanilla transformer-based LLMs by replacing the MLP module with a set of smaller MLP modules (experts) and sparsely activating partial amount of experts for each input token [17]. Despite its performance benefits, the deployment of MoE-based models often incur additional memory overhead to maintain the non-activated experts in memory, which is valuable but limited for existing accelerators such as GPU and TPU [16].
@@
While diverse structured pruning methods have been proposed to reduce deployment memory costs by removing different model components while minimizing the performance degradation [32, 12, 34], we observe that these approaches often exhibit suboptimal performance and unstable degradation when applied to MoE models. Specifically, we identify three critical dimensions where existing methods fall short: model architectures, calibration data sources and calibration sample sizes, as shown by experiments in Section 5.3. These limitations are evident across two main categories of structured pruning approaches for MoE models: general structured pruning and expert pruning as shown in Figure 1 (a). First, general structured pruning methods that remove model layers (Angular [11]) or weight matrix channels (FLAP [2]) fail to account for the sparse computation scheme of MoE models when evaluating component importance, resulting in inconsistent performance across the aforementioned three dimensions. Second, existing expert pruning methods such as MC-SMoE [18] and RS [12] remove experts from MoE models primarily based on the expert access frequency. However, as shown in Figure 1 (b), this feature alone fails to fully capture the expert redundancy. Besides, these methods lack mechanisms to recover the performance loss caused by pruning.
@@
Figure 1: (a) Different structured pruning methods. (b) Layer-wise normalized expert access frequency and output variance of Deepseek-V2-Lite for three downstream tasks. Experts with high access frequency or output variances are the same across downstream tasks. Expert in blue circles has both high frequency and variance. Expert in red circles only has high variance. Expert in green circles only has high frequency. Similar observations on other models and tasks are in Appendix B.
@@
To improve the effectiveness and robustness of structured pruning for MoE models, this paper proposes a novel expert pruning method, Mixture-of-Novices-and-Experts (MoNE) which replaces redundant experts with a lightweight structure, novice. Specifically, to prune an MoE model with MoNE, it first evaluates the expert redundancy by the access frequency and the output variance for each expert on a calibration dataset. Then, it identifies and removes redundant experts that show low access frequency and stable output activations. Finally, the unbiased estimation of the expert output is employed as the lightweight novice to reduce the computation and memory overhead for the pruned expert. The intuition behind MoNE is that experts with low access frequency contribute less to the final outputs and experts whose outputs have low variance can be replaced with a constant but introduce less discrepancy. Moreover, Figure 1 (b) reveals that experts with less redundancy identified by MoNE exhibit strong consistency across various downstream tasks.
@@
  • We propose a novel expert pruning method named MoNE which replaces redundant experts with lightweight novices to compress MoE models.
@@
  • Extensive experiment results demonstrate that MoNE consistently outperforms baseline methods under varying MoE architectures, calibration data sources and calibration sample sizes. Notably, it improves the average zero shot accuracy across nine downstream tasks by up to 2.71 with 25% pruning ratio and 3.61 with 50% pruning ratio.
@@ 2 Related Work
Model pruning compresses a model by removing certain redundant model parameters while preserving accuracy. Existing pruning methods generally fall into two categories: unstructured pruning and structured pruning. Unstructured pruning eliminates any model parameter that has minimal impact on model performance. Methods such as SparseGPT [9], Wanda [30], SparseLLM [3] excel in maintaining accuracy while achieving high compression ratios. However, the resulting irregular sparsity patterns hinder efficient representation and execution on hardware accelerators.
In contrast, structured pruning removes certain modules of a model, preserving hardware-friendly structures. Early researches prune redundant transformer layers of a LLM [8, 19, 11]. LLM-Pruner [23] and FLAP [2] remove rows or columns of individual weight matrices. Recent work also proposes to delete components such as attention, MLP or MoE modules within each transformer layer [32, 12]. Minitron [26] and Sheared LLaMA [34] combine different granularity and automatically search for the optimal structures to prune. Despite their versatility, existing structured pruning methods often exhibit inconsistent performance across MoE architectures.
This work focuses on expert pruning, a unique direction of structured pruning for MoE models [17, 20]. Expert pruning targets on deleting individual experts for each layer to compress an MoE model. Previous expert pruning methods either require exhaustive search to identify redundant experts [22], or heavily rely on retraining to recover accuracy due to the suboptimal pruning performance [18, 12]. However, the exhaustive search is not applicable to modern MoE model architectures such as Deepseek [7, 21] or OLMoE [25], as their MoE layer contains 64 experts or even more, yielding a tremendous search space that is intractable. Retraining obscures the advantages of expert pruning over other structured pruning methods.
@@ 3.1 Mixture-of-Experts (MoE)
MoE-based LLMs replace the traditional MLP module in the transformer layer with MoE module. Each MoE module consists of a router network $G$ and a set of experts =$\mathcal{E}=\{E_{1},E_{2},\dots,E_{M}\}$, where $M$ is the number of experts and each expert is a smaller MLP. Let $\textbf{x}\in\mathbb{R}^{d}$ be the hidden state of an input token, where $d$ is the hidden size of the model, the output of an MoE module is computed as:
@@
The output of the router network $G(\textbf{x})\in\mathbb{R}^{M}$ represents the routing scores for all experts, and $\mathcal{S}_{k,\textbf{x}}\subseteq\mathcal{E}$ denotes the top $k$ experts with the highest routing scores for input x. The final output of the MoE module is the weighted sum of outputs from the top $k$ experts. While Equation 1 captures the general MoE computation, implementations for $G$ and $E_{i}$ may vary across model architectures [7, 20, 25, 21].
@@ 3.2 Expert pruning formulation
Previous studies have revealed that not all experts contribute equally, and pruning less important ones can reduce memory overhead with marginal performance degradation [22, 18, 14]. However, searching for the target experts to prune at the global model perspective falls into a tremendous search space, as the number of experts per transformer layer increases with the evolving of the MoE model architectures [17, 15, 7, 20]. Following the layer-wise pruning scheme [9, 2, 22, 19], our goal of expert pruning is to identify a subset of redundant experts $\mathcal{P}\subseteq\mathcal{E}$ such that we can minimize the output difference after compressing their parameters:
@@
To achieve this goal, the core problem is twofold: (1) find a metric to evaluate the importance of the experts in each layer, so that we can identify the expert subset $\mathcal{P}$, and (2) find an pruning method to compress the parameters of $\mathcal{P}$, so that we can reduce the model size. While $\mathcal{E}\setminus\mathcal{P}$ implies directly removing redundant experts [12, 22], existing methods have also explored expert merging to mitigate the expert redundancy [18].
@@ 4.1 MoNE framework
MoE models are often trained with auxiliary losses to ensure load balance among experts in each layer, enabling each expert to learn certain aspect of knowledge [17, 7, 25]. However, most existing expert pruning methods directly remove experts [12, 18], often leading to inconsistent performance drops across different model architectures or calibration data. MoNE addresses this issue by introducing lightweight structures called novices to replace the pruned experts. A novice is designed to capture the essential knowledge previously held by the removed experts. In contrast to simply removing redundant experts, MoNE compensates for knowledge loss by leveraging novices, thereby preserving the overall performance of the model while maintaining compression efficiency. Specifically, the output of the MoNE is computed as:
@@ 4.2 Expert redundancy evaluation
The routing scores and access frequencies of the router network $G$ serve as strong indicators of the overall redundancy of an expert [12, 18]. Intuitively, experts which are rarely selected or consistently assigned lower routing scores are likely to have a minimal impact on the model’s output. As shown in Figure 1 (b), we can identify typical experts in blue and green circles that show consistent high frequency over the three downstream tasks. Notably, the expert in green circles only has high frequency. Therefore, the frequency and variance information can complement the discrepancy ignored by each other. Based on this observation, we define the frequency-based redundancy $\phi_{i}^{freq}$ of the expert $E_{i}$ as the average routing score across a calibration dataset $\mathcal{C}$ of which $E_{i}$ is among the top $k$ selected experts. Formally, the frequency-based redundancy $\phi_{i}^{freq}$ is defined as:
@@ 4.3 Expert replacement with novice
$\underset{E_{i}\in\mathcal{P}}{\min}\|\sum_{\textbf{x}\in\mathcal{C}}(E_{i}(\textbf{x})-N_{i})\|_{2}$ (8)
@@ 5.1 Experiment setup
To validate the effectiveness and robustness of MoNE, we conducted structured pruning on three open source MoE models with diverse architectures and model scales: OLMoE [25], Moonlight [21] and DeepSeek-V2-Lite [7]. OLMoE has 7B parameters with 1B activated parameters per token. Both Moonlight and Deepseek-V2-Lite have 16B parameters with 3B activated pamaters per token. OLMoE and Moonlight represent SOTA MoE models at their respective scales. We chose the base version of the three models for experiments [^2] OLMoE: https://huggingface.co/allenai/OLMoE-1B-7B-0125, Moonlight: https://huggingface.co/moonshotai/Moonlight-16B-A3B, Deepseek V2 Lite: https://huggingface.co/deepseek-ai/DeepSeek-V2-Lite.
@@
MoNE 42.32 64.81 67.19 85.00 40.13 40.80 78.07 64.62 66.46 61.04
FLAP 29.18 54.92 62.17 68.00 30.51 29.60 67.57 55.23 56.27 50.39
MC-SMoE 24.49 31.44 59.33 67.00 23.01 26.00 53.92 53.43 53.12 43.53
RS 21.50 28.62 39.45 61.00 23.27 26.00 52.34 51.99 51.85 39.56
MoNE 28.16 40.24 63.12 78.00 25.21 32.40 63.33 60.65 63.54 50.52
@@
MoNE 55.89 80.60 79.57 90.00 55.23 46.80 80.85 61.01 71.98 69.10
FLAP 33.87 61.36 63.30 75.00 36.80 36.00 69.37 57.04 62.12 54.98
MC-SMoE 29.52 47.94 59.54 79.00 23.94 31.40 67.30 57.04 60.46 50.68
RS 37.80 58.42 70.86 89.00 23.27 38.00 78.18 57.76 70.80 58.23
MoNE 43.09 70.03 76.12 90.00 23.57 40.80 78.78 58.84 70.17 61.27
@@
MoNE 46.67 74.62 78.47 90.00 49.05 43.00 79.76 62.09 71.43 66.12
FLAP 35.24 60.31 69.66 79.00 36.13 35.20 74.76 56.68 64.09 56.79
MC-SMoE 24.57 35.82 56.36 67.00 28.29 27.40 61.37 52.35 53.91 45.23
RS 36.01 57.45 57.98 89.00 24.91 40.80 78.02 54.15 62.75 55.67
MoNE 37.20 67.17 73.39 84.00 42.30 36.80 75.30 59.57 67.88 60.40
Baseline methods
We selected structured pruning methods for different structures as baseline. Notably, unless explicitly stated, we did not apply any weight update to compare the effect of pruning methods. Specifically, for general structured pruning methods, we used Angular for layer pruning [11], which evaluates the layer importance by the angular distance between the input activations for different layers, and we used FLAP for weight pruning [2], which evaluates the channel importance by the fluctuation of the input activations and compensates the performance loss with the averaged output activations. For expert pruning methods, we adopted the expert merging method in MC-SMoE [18] for one of the expert pruning baselines. Another expert pruning baseline is RS [12], which uses routing scores to evaluate the expert importance and discards less accessed ones. As there are 64 experts and 8 activated experts per layer for the three MoE models, the two expert pruning methods will not change the activated parameters. In contrast, the proposed MoNE adaptively reduces the activated parameters for tokens routed to novices.
Implementation details
We tested two pruning ratios: 25% and 50%. To demonstrate the robustness of MoNE to calibration data, we conducted experiments on two calibration data sources: Zyda2 [31] and C4 [27]. Both datasets are constructed for LLM pretraining and C4 is commonly used as the calibration dataset for model pruning [19, 9, 11, 34]. Besides, we also investigated the performance under three calibration sample sizes: 100, 500 and 1000 in Section 5.3.
To evaluate performance recovery capabilities, we conducted continued pretraining on the 25% compressed OLMoE model pruned by 100 Zyda2 samples, as only this model releases its pretraining dataset, OLMoE-mix-0924333https://huggingface.co/datasets/allenai/OLMoE-mix-0924. The sequence length was set to 4096 and the global token size per step was 4M. Each pruned model was trained with 2B tokens, i.e., 512 steps, and the peak and minimum learning rate (lr) were 5e-5 and 5e-6, respectively. We employed the cosine lr scheduler with 50 warm up steps. Other hyperparameters were the same as the original configuration for OLMoE [25]. All the experiments could run on a single H20 GPU, but we accelerated the training with 16 H20 GPUs.
Following previous researches [23, 3, 19, 34, 2], we adopted lm-evaluation-harness [^4] https://github.com/EleutherAI/lm-evaluation-harness [10] to measure the zero shot accuracy and average results on nine downstream tasks: Arc-c and Arc-e [6], BoolQ [5], COPA [28], MMLU [13], OBQA [24], PIQA [4], RTE [33] and Winogrande [29].

### v2

@@ Abstract
Mixture-of-Experts (MoE) enables efficient scaling of large language models by activating only a subset of experts per input token. However, deploying MoE-based models incurs significant memory overhead due to the need to retain all experts in memory. While structured pruning is promising to reduce memory costs, existing methods often show suboptimal performance and unstable degradation in three dimensions: model architectures, calibration data sources, and calibration sample sizes. This paper proposes Mixture-of-Novices-and-Experts (MoNE), a novel expert pruning method that replaces redundant experts with lightweight novices to achieve effective and robust model compression. MoNE evaluates expert redundancy based on two metrics: access frequency and output variance. Experts exhibiting low usage and stable outputs are pruned and replaced with lightweight novices—unbiased estimations of their original outputs—minimizing performance degradation. Extensive experiments demonstrate that MoNE consistently outperforms baseline methods with minimal accuracy degradation across the three dimensions, confirming its effectiveness and robustness. Notably, it outperforms baselines by up to 2.72 for the average zero shot accuracy across nine downstream tasks under 25% pruning ratio, with only 0.14 performance drop for Qwen2-57B-A14B. The code is available at https://github.com/zxgx/mode-pd.
@@ 1 Introduction
Mixture-of-Experts (MoE) has emerged as a powerful architecture for advancing the capabilities of large language models (LLMs) (Liu et al., 2024; 2025; Muennighoff et al., 2025). MoE-based LLMs achieve higher parameter efficiency than vanilla transformer-based LLMs by replacing the MLP module with a set of smaller MLP modules (experts) and sparsely activating partial experts for each input token (Lepikhin et al., 2021). Despite its performance benefits, the deployment of MoE-based models often incur additional memory overhead to maintain the non-activated experts in memory, which is valuable but limited for existing accelerators such as GPU and TPU (Jouppi et al., 2023).
@@
While diverse structured pruning methods have been proposed to reduce deployment memory costs by removing different model components while minimizing the performance degradation (Voita et al., 2019; He et al., 2024; Xia et al., 2024; Zhao et al., 2026; 2025a; 2025b), we observe that these approaches often exhibit suboptimal performance and unstable degradation when applied to different MoE models. Specifically, we identify three critical dimensions where existing methods fall short: model architectures, calibration data sources and calibration sample sizes, as shown by experiments in Section 5.3. These limitations are evident across two main categories of structured pruning approaches for MoE models: general structured pruning and expert pruning as shown in Figure 1 (a). First, general structured pruning methods that remove model layers (Angular (Gromov et al., 2025)) or weight matrix channels (FLAP (An et al., 2024)) fail to account for the sparse computation scheme of MoE models when evaluating model component importance, resulting in inconsistent performance drop across the aforementioned three dimensions. Second, existing expert pruning methods such as MC-SMoE (Li et al., 2024) and RS (He et al., 2024) remove experts from MoE models primarily based on the expert access frequency. However, as shown in Figure 1 (b), this feature alone fails to fully capture the expert redundancy. Besides, these methods lack mechanisms to recover the performance loss caused by pruning.
@@
Figure 1: (a) Different structured pruning methods. (b) Layer-wise normalized expert access frequency and output variance of Deepseek-V2-Lite for three downstream tasks. Experts with high access frequency or output variances are the same across downstream tasks. Expert in blue circles has both high frequency and variance. Expert in red circles only has high variance. Expert in green circles only has high frequency. Similar observations on other models and tasks are in Appendix D.
@@
To improve the effectiveness and robustness of structured pruning for MoE models, this paper proposes a novel expert pruning method, Mixture-of-Novices-and-Experts (MoNE) which replaces redundant experts with a lightweight structure, novice. Specifically, to prune an MoE model with MoNE, it first evaluates the expert redundancy by the access frequency and the output variance for each expert on a calibration dataset. Then, it identifies and prunes redundant experts that show low access frequency and stable output activations to reduce the memory overhead from redundant experts. Finally, the unbiased estimation of the pruned expert output is employed as the lightweight novice to reclaim the performance loss caused by the pruned expert. The intuition behind MoNE is that experts with low access frequency contribute less to the final outputs and experts whose outputs have low variance can be replaced with a constant but introduce less discrepancy. Moreover, Figure 1 (b) reveals that experts with less redundancy identified by MoNE exhibit strong consistency across various downstream tasks.
@@
  • We propose a novel expert pruning method named MoNE which replaces redundant experts with lightweight novices to compress MoE models with minimal performance loss.
@@
  • Extensive experiment results demonstrate that MoNE consistently outperforms baseline methods under varying MoE architectures, calibration data sources and calibration sample sizes. Notably, it outperforms baselines by up to 2.72 for the average zero shot accuracy across nine downstream tasks under 25% pruning ratio, with only 0.14 performance drop for Qwen2-57B-A14B.
@@ 2 Related Work
Model pruning compresses a model by removing certain redundant model parameters while preserving accuracy. Existing pruning methods generally fall into two categories: unstructured pruning and structured pruning. Unstructured pruning eliminates any model parameter that has minimal impact on model performance. Methods such as SparseGPT (Frantar and Alistarh, 2023), Wanda (Sun et al., 2024), SparseLLM (Bai et al., 2024) excel in maintaining accuracy while achieving high compression ratios. However, the resulting irregular sparsity patterns hinder efficient representation and execution on hardware accelerators.
In contrast, structured pruning removes certain modules of a model, preserving hardware-friendly structures. Early researches prune redundant transformer layers of a LLM (Fan et al., 2020; Ling et al., 2024; Gromov et al., 2025). LLM-Pruner (Ma et al., 2023), FLAP (An et al., 2024), MoE-Pruner (Xie et al., 2024) and SlimMoE (Li et al., 2025) remove rows or columns of individual weight matrices. Recent work also proposes to delete components such as attention, MLP or MoE modules within each transformer layer (Voita et al., 2019; He et al., 2024). Minitron (Muralidharan et al., 2024) and Sheared LLaMA (Xia et al., 2024) combine different granularity and automatically search for the optimal structures to prune. Despite their versatility, existing structured pruning methods often exhibit inconsistent performance across MoE architectures.
This work focuses on expert pruning, a unique direction of structured pruning for MoE models (Lepikhin et al., 2021; Liu et al., 2024). Expert pruning targets on deleting individual experts for each layer to compress an MoE model. Previous expert pruning methods either require exhaustive search to identify redundant experts (Lu et al., 2024), or heavily rely on retraining to recover accuracy due to the suboptimal pruning performance (Li et al., 2024; He et al., 2024). However, the exhaustive search is not applicable to modern MoE model architectures such as Deepseek (DeepSeek-AI, 2024; Liu et al., 2025), OLMoE (Muennighoff et al., 2025) or Qwen (Qwen, 2024; Yang et al., 2025), as their MoE layer contains 64 experts or even more, yielding a tremendous search space that is intractable. Retraining obscures the advantages of expert pruning over other structured pruning methods.
@@ 3.1 Mixture-of-Experts (MoE)
MoE-based LLMs replace the traditional MLP module in the transformer layer with MoE module. Each MoE module consists of a router network $G$ and a set of experts $\mathcal{E}=\{E_{1},E_{2},\dots,E_{M}\}$, where $M$ is the number of experts and each expert is a smaller MLP. Let $\textbf{x}\in\mathbb{R}^{d}$ be the hidden state of an input token, where $d$ is the hidden size of the model, the output of an MoE module is computed as:
@@
The output of the router network $G(\textbf{x})\in\mathbb{R}^{M}$ represents the routing scores for all experts, and $\mathcal{S}_{k,\textbf{x}}\subseteq\mathcal{E}$ denotes the top $k$ experts with the highest routing scores for input x. The final output of the MoE module is the weighted sum of outputs from the top $k$ experts. While Equation 1 captures the general MoE computation, implementations for $G$ and $E_{i}$ may vary across model architectures (DeepSeek-AI, 2024; Liu et al., 2024; Muennighoff et al., 2025; Liu et al., 2025).
@@ 3.2 Expert pruning formulation
Previous studies have revealed that not all experts contribute equally, and pruning less important ones can reduce memory overhead with marginal performance degradation (Lu et al., 2024; Li et al., 2024; Huang et al., 2025). However, searching for the target experts to prune at the global model perspective falls into a tremendous search space, as the number of experts per transformer layer increases with the evolving of the MoE model architectures (Lepikhin et al., 2021; Jiang et al., 2024; DeepSeek-AI, 2024; Liu et al., 2024). Following the layer-wise pruning scheme (Frantar and Alistarh, 2023; An et al., 2024; Lu et al., 2024; Ling et al., 2024), our goal of expert pruning is to identify a subset of redundant experts $\mathcal{P}\subseteq\mathcal{E}$ such that we can minimize the output difference after compressing their parameters:
@@
To achieve this goal, the core problem is twofold: (1) find a metric to evaluate the importance of the experts in each layer, so that we can identify the expert subset $\mathcal{P}$, and (2) find an pruning method to compress the parameters of $\mathcal{P}$, so that we can reduce the model size. While $\mathcal{E}\setminus\mathcal{P}$ implies directly removing redundant experts (He et al., 2024; Lu et al., 2024), existing methods have also explored expert merging to mitigate the expert redundancy (Li et al., 2024).
@@ 4.1 MoNE framework
MoE models are often trained with auxiliary losses to ensure load balance among experts in each layer, enabling each expert to learn certain aspect of knowledge (Lepikhin et al., 2021; DeepSeek-AI, 2024; Muennighoff et al., 2025). However, most existing expert pruning methods directly remove experts (He et al., 2024; Li et al., 2024), often leading to inconsistent performance drops across different model architectures or calibration data. MoNE addresses this issue by introducing lightweight structures called novices to replace the pruned experts. A novice is designed to capture the essential knowledge previously held by the removed experts. In contrast to simply removing redundant experts, MoNE compensates for knowledge loss by leveraging novices, thereby preserving the overall performance of the model while maintaining compression efficiency. Specifically, the output of the MoNE is computed as:
@@ 4.2 Expert redundancy evaluation
The routing scores and access frequencies of the router network $G$ serve as strong indicators of the overall redundancy of an expert (He et al., 2024; Li et al., 2024). Intuitively, experts which are rarely selected or consistently assigned lower routing scores are likely to have a minimal impact on the model’s output. As shown in Figure 1 (b), we can identify typical experts in blue and green circles that show consistent high frequency over the three downstream tasks. Notably, the expert in green circles only has high frequency. Therefore, the frequency and variance information can complement the discrepancy ignored by each other. Based on this observation, we define the frequency-based redundancy $\phi_{i}^{freq}$ of the expert $E_{i}$ as the average routing score across a calibration dataset $\mathcal{C}$ of which $E_{i}$ is among the top $k$ selected experts. Formally, the frequency-based redundancy $\phi_{i}^{freq}$ is defined as:
@@ 4.3 Expert replacement with novice
$\underset{E_{i}\in\mathcal{P}}{\min}\sum_{\textbf{x}\in\mathcal{C}}(\|E_{i}(\textbf{x})-N_{i}\|_{2})$ (8)
@@ 5.1 Experiment setup
To validate the effectiveness and robustness of MoNE, we conducted structured pruning on five open source MoE models with diverse architectures and model scales: OLMoE (Muennighoff et al., 2025), Moonlight (Liu et al., 2025), DeepSeek-V2-Lite (DeepSeek-AI, 2024), Qwen2-57B-A14B (Qwen, 2024) and Qwen3-30B-A3B (Yang et al., 2025). OLMoE has 7B parameters with 1B activated parameters per token. Both Moonlight and Deepseek-V2-Lite have 16B parameters with 3B activated pamaters per token. OLMoE and Moonlight represent SOTA MoE models at their respective scales. To demonstrate scalability to larger architectures, we also consider the Qwen series: Qwen3-30B-A3B with 30B parameters and 3B activated per token, and Qwen2-57B-A14B with 57B parameters and 14B activated per token. We chose the base version of the five models for experiments.
Baseline methods
We selected structured pruning methods for different structures as baseline. Notably, unless explicitly stated, we did not apply any weight update to compare the effect of pruning methods. Specifically, for general structured pruning methods, we used Angular for layer pruning (Gromov et al., 2025), which evaluates the layer importance by the angular distance between the input activations for different layers, and we used FLAP for weight pruning (An et al., 2024), which evaluates the channel importance by the fluctuation of the input activations and compensates the performance loss with the averaged output activations. For expert pruning methods, we adopted the expert merging method in MC-SMoE (Li et al., 2024) for one of the expert pruning baselines. Another expert pruning baseline is RS (He et al., 2024), which uses routing scores to evaluate the expert importance and discards less accessed ones.
Implementation details
We tested two pruning ratios: 25% and 50%. To demonstrate the robustness of MoNE to calibration data, we conducted experiments on two calibration data sources: Zyda2 (Tokpanov et al., 2024) and C4 (Raffel et al., 2020). Both datasets are constructed for LLM pretraining and C4 is commonly used as the calibration dataset for model compression (Ling et al., 2024; Frantar and Alistarh, 2023; Gromov et al., 2025; Xia et al., 2024). Besides, we also investigated the performance under three calibration sample sizes: 100, 500 and 1000 in Section 5.3.
Evaluation protocol
Following previous researches (Ma et al., 2023; Bai et al., 2024; Ling et al., 2024; Xia et al., 2024; An et al., 2024), we adopted lm-evaluation-harness [^1] https://github.com/EleutherAI/lm-evaluation-harness (Gao et al., 2024) to measure the zero shot accuracy and average results on nine downstream tasks: Arc-c and Arc-e (Clark et al., 2018), BoolQ (Clark et al., 2019), COPA (Roemmele et al., 2011), MMLU (Hendrycks et al., 2021), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), RTE (Wang et al., 2019) and Winogrande (Sakaguchi et al., 2021).
Though more complex downstream tasks such as coding (Zhang et al., 2025), math (Lightman et al., 2023) or reasoning (Lin et al., 2025) exist, these tasks are still challenging for full LLMs (Yang et al., 2025; Liu et al., 2024). Moreover, existing study shows that model compression for complex tasks often requires additional task-specific fine-tuning (Sarkar et al., 2024; Chen et al., 2022), which is beyond the scope of this work and we consider it a promising direction of future work.
5.2 Effectiveness evaluation
@@
MoNE (Ours) 42.32 64.81 67.19 85.00 40.13 40.80 78.07 64.62 66.46 61.04
@@
MoNE (Ours) 55.89 80.60 79.57 90.00 55.23 46.80 80.85 61.01 71.98 69.10
@@
MoNE (Ours) 46.67 74.62 78.47 90.00 49.05 43.00 79.76 62.09 71.43 66.12
0% Qwen2-57B-A14B 49.66 69.44 86.45 93.00 74.06 44.20 81.23 74.73 74.27 71.89
25% Angular 29.44 54.17 59.51 70.00 23.92 32.80 70.02 54.87 49.57 49.37
FLAP 50.00 72.85 86.91 91.00 65.02 45.40 81.12 77.62 76.09 71.78
MC-SMoE 46.67 66.25 86.45 88.00 69.46 43.40 80.20 74.73 75.14 70.03
RS 49.15 69.78 84.77 87.00 70.99 44.80 81.34 74.37 74.19 70.71
MoNE (Ours) 49.66 68.73 86.88 94.00 71.64 45.20 81.07 75.45 73.09 71.75
0% Qwen3-30B-A3B 55.89 79.42 88.69 84.00 77.82 44.80 80.30 82.31 70.64 73.76
25% Angular 44.62 68.60 80.52 77.00 59.75 40.40 75.30 70.40 62.51 64.34
FLAP 50.85 76.68 85.72 85.00 69.43 42.80 77.31 81.95 70.17 71.10
MC-SMoE 52.73 76.98 88.75 83.00 72.25 44.40 79.71 80.87 70.40 72.12
RS 53.75 78.32 88.53 85.00 74.60 43.00 79.92 79.78 69.61 72.50
MoNE (Ours) 56.14 79.17 89.11 85.00 74.04 43.00 79.27 77.98 70.48 72.69
Table 2: Maximum pruning ratios with 1% accuracy loss after MoNE pruning using 100 calibration samples from Zyda-2 dataset.
Model Max. pruning ratio Avg. perf before pruning Avg. perf after pruning
OLMoE 16% 66.63 66.00
Moonlight 16% 71.56 70.59
Deepseek-V2-Lite 20% 67.16 66.31
Qwen2-57B-A14B 25% 71.89 71.75
Qwen3-30B-A3B 24% 73.76 73.61
This section validates the effectiveness of MoNE by comparing the zero shot performance of 25% pruned models with 100 calibration samples from the Zyda2 dataset. The results are presented in Table 1 and Table 2. Results under 50% pruning ratio are extended to Table 10 in Appendix C.
Table 1 indicates that MoNE consistently outperforms baseline methods in terms of the average accuracy on the nine tasks. In particular, it shows average accuracy improvement as large as 2.72 for the pruned OLMoE compared to baseline methods, and it incurs accuracy drop as small as only 0.14 for the pruned Qwen2-57B-A14B. Furthermore, MoNE-pruned models can achieve either the best or the second best result for individual tasks under most settings.
An interesting observation is that all the three expert pruning methods, MC-SMoE, RS and MoNE can achieve results on par with or even better than the original models on certain tasks. The specific examples are shown in green background in Table 1. All these results indicate that there is indeed redundancy existing in the expert level for the examined MoE models, and expert pruning can rule out such redundancy to achieve even better results on these tasks. Among the three expert pruning methods, MoNE consistently surpasses the two baseline methods, demonstrating its strong capability.
Table 2 evaluates the maximum pruning ratios that maintain accuracy loss within 1% across different MoE models. As shown in Table 2, larger and more powerful models such as Qwen2-57B-A14B and Qwen3-30B-A3B tolerate more aggressive pruning (25% and 24%, respectively) compared to smaller models like OLMoE, Moonlight, and Deepseek-V2-Lite (16–20%). This finding highlights the scalability of proposed MoNE method for large model scales. In addition, combining the observations from Table 1 and Table 2, we argue that larger MoE models may have stronger capability in language modeling but also contain increasing parameter redundancy at the expert level, and MoNE can efficiently eliminates such redundancy at minimal performance degradation.
5.3 Robustness evaluation